OpenStudy (anonymous):

Identify the first step using the elimination method to solve the following system of equations: 6x + 3y = 16 7x - 7y = 14 A. 7(6x + 3y = 12) 6(7x - 7y = 14) B. 7(6x + 3y = 12) 3(7x - 7y = 14) C. 7(6x + 3y = 12) -3(7x - 7y = 14) D. -14(6x + 3y = 12) 16(7x - 7y = 14)

OpenStudy (anonymous):

\[WELCOME'TO'OPEN'STUDY!!!!! ")\]

OpenStudy (anonymous):

(2x + y) = 12 7(x - y) = 14 2x + y = 4 x - y = 2 3x = 6 x = 2 These'' are'' the''steps \to use elimination\]